OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 4 and a6 = 39 ?

OpenStudy (muhammad_nauman_umair):

a6=a1+5d here d=a6-a1/5 as a6=39 and a1=1 d=38/5 and a32=a1+31d a32=1+31(38/5)

OpenStudy (anonymous):

A) 209 B) 213 C) 217 D) 221 those are the possible answers

OpenStudy (callisto):

Hmm.. the above procedures are correct, but not the number. a6=a1+5d d=a6-a1/5, where a6 = 39 and a1=4 d=(39-4)/5 = 7 a32=a1+(32-1)d a32=4+31(7) = Can you work it out now?

OpenStudy (anonymous):

yup 221 correct..

OpenStudy (muhammad_nauman_umair):

sorry ... for my mistake

OpenStudy (campbell_st):

1st find d the common difference use \[T _{n} = a + (n - 1)d\] n = 6 T6 = 39 and a = 4 \[39 = 4 + (6 - 1)d\] \[35 = 5d\] d = 7 to find the 32 term, use the above formula with a = 4 d = 7 and n = 32 then just evaluate it